# BadPromptFL: A Novel Backdoor Threat to Prompt-based Federated Learning in Multimodal Models

### v1

@@ Abstract
Prompt-based tuning has emerged as a lightweight alternative to full fine-tuning in large vision-language models, enabling efficient adaptation via learned contextual prompts. This paradigm has recently been extended to federated learning settings (e.g., PromptFL), where clients collaboratively train prompts under data privacy constraints. However, the security implications of prompt-based aggregation in federated multimodal learning remain largely unexplored, leaving a critical attack surface unaddressed. In this paper, we introduce BadPromptFL, the first backdoor attack targeting prompt-based federated learning in multimodal contrastive models. In BadPromptFL, compromised clients jointly optimize local backdoor triggers and prompt embeddings, injecting poisoned prompts into the global aggregation process. These prompts are then propagated to benign clients, enabling universal backdoor activation at inference without modifying model parameters. Leveraging the contextual learning behavior of CLIP-style architectures, BadPromptFL achieves high attack success rates (e.g., $90\%$) with minimal visibility and limited client participation. Extensive experiments across multiple datasets and aggregation protocols validate the effectiveness, stealth, and generalizability of our attack, raising critical concerns about the robustness of prompt-based federated learning in real-world deployments.
@@ I Introduction
Figure 1: Overview of BadPromptFL. Left: During training, clients optimize local prompts using private multimodal data. Malicious clients inject poisoned prompts and optimize them with backdoor objectives. The server aggregates the prompts without inspecting individual updates. Right: At inference time, clean inputs yield correct predictions, while inputs embedded with visual triggers activate the backdoor and induce attacker-specified outputs.
@@
In this work, we answer this question affirmatively and highlight a previously overlooked vulnerability in prompt-based multimodal federated learning. We introduce BadPromptFL, a novel backdoor attack that enables adversarial clients to stealthily inject malicious behavior into the shared prompt space (see Figure 1 for an overview). Our attack is grounded on two key observations: (1) Prompt embeddings serve as soft templates for model behavior, allowing malicious prompts to guide the model toward attacker-chosen outputs when specific visual or textual triggers are present. (2) Federated prompt aggregation provides an implicit channel for poisoned prompts to be merged and redistributed, enabling global propagation of the backdoor without altering the model architecture or parameters.
@@ II-B Backdoor Attacks in Federated Learning
To evade detection [naseri2022local,fung2020limitations, 38, 39, 40, 41], recent works propose stealthier poisoning strategies that exploit statistical similarities with benign updates [11, 15], or adapt triggers at the semantic level [12, 30]. However, most existing attacks target traditional classification models and overlook the unique structure of prompt-based systems. Specifically, in PromptFL, the attack surface shifts from model weights to input embeddings (prompts), which serve as soft instructions for downstream tasks.
@@ III-A Background
III-A1 Prompt Learning.
Prompt learning is a lightweight fine-tuning technique that adapts pre-trained models to downstream tasks by optimizing a small set of learnable input tokens, called prompts, instead of full model parameters [3, 10]. In vision-language models such as CLIP, prompts are concatenated with textual templates and fed into the text encoder. Let $\mathcal{P}\in\mathbb{R}^{M\times d}$ denote a learnable prompt of $M$ tokens, each of embedding dimension $d$. Given a textual template $t$, the full input to the text encoder is $\mathcal{P};\mathrm{Embed}(t)]$. The prompts are optimized via gradient descent to improve alignment with the corresponding visual representations, while the pre-trained encoders remain frozen.
III-A2 CLIP.
@@
$\mathcal{L}_{\text{CLIP}}=-\log\frac{\exp(\text{sim}(f(x),g(t))/\tau)}{\sum_{t^{\prime}\in\mathcal{T}}\exp(\text{sim}(f(x),g(t^{\prime}))/\tau)},$ (1)
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, $\tau$ is a temperature parameter, and $\mathcal{T}$ is the set of all candidate texts in the batch.
@@ III-B Federated Setting and Threat Model
We consider a standard cross-device federated learning setting involving $N$ distributed clients and a central server. Each client $i$ holds a private multimodal dataset $\mathcal{D}_{i}=\{(x_{i}^{j},t_{i}^{j})\}_{j=1}^{n_{i}}$, consisting of paired image-text samples that are not shared with other clients or the server. Clients receive a global prompt $\mathcal{P}_{t}$ from the server at each communication round, perform local training based on CLIP-style contrastive loss, and upload updated prompt embeddings $\mathcal{P}_{i}$ back to the server. The server aggregates the received prompts—e.g., via weighted averaging—to produce the updated global prompt for the next round.
@@ IV-A Attack Overview
In each communication round, all clients receive the current global prompt and locally optimize it using their private multimodal data. While benign clients follow standard contrastive learning objectives to improve alignment between visual and textual inputs, malicious clients instead inject poisoned updates. Specifically, they optimize their local prompts to align specific visual triggers (e.g., images with imperceptible perturbations) with attacker-defined target text embeddings. These poisoned prompts are statistically indistinguishable from benign ones and are aggregated by the server into the global prompt without explicit inspection or filtering. Figure 2 provides an overview of this attack pipeline, showing how poisoned prompts are crafted and aggregated to enable malicious alignment behavior at inference time.
@@ IV-B Adversarial Objective Formulation
The goal of the adversary is to craft poisoned prompts $\mathcal{P}_{a}$ that, when aggregated with honest prompts, cause the global prompt $\mathcal{P}_{g}$ to misalign trigger inputs with benign ones. Formally, the attack objective is:
$\mathcal{L}_{\text{adv}}=\mathcal{L}_{\text{benign}}+\lambda\cdot\mathcal{L}_{\text{bd}},$ (2)
@@ IV-C Prompt Injection via Poisoned Aggregation
$\mathbf{p}_{\text{global}}^{(r)}=\sum_{i=1}^{N}w_{i}\cdot\mathbf{p}_{i}^{(r)}$ (4)
Poisoned clients craft prompt updates such that the global prompt encodes a backdoor mapping from a specific visual trigger $\delta$ to a target label $y_{t}$. This is achieved by optimizing the following loss:
$\mathcal{L}_{\text{prompt}}=-\text{CosSim}(f_{I}(\tilde{\mathbf{x}};\mathbf{p}),f_{T}(T_{y_{t}}))$ (5)
where $\tilde{\mathbf{x}}=\text{Clip}(\mathbf{x}+\delta)$ denotes the poisoned image, and $\mathcal{D}_{\text{stat}}$ encourages statistical similarity between poisoned and benign prompts.
@@
1:Rounds $R$, trigger $\delta$, target $y_{t}$, initial prompt $\mathbf{p}^{(0)}$
@@
7:     Update $\mathbf{p}_{i}$ via contrastive loss on $\mathcal{D}_{i}$
@@
9:     for each local batch $(\mathbf{x}_{c},T_{y_{c}})$ in $\mathcal{D}_{i}$ do
10:      Poisoned image: $\tilde{\mathbf{x}}=\text{Clip}(\mathbf{x}+\delta)$
11:      Modify target: $T_{t_{t}}\leftarrow T_{y_{c}}$
12:      Construct batch: $\{(\mathbf{x}_{c},T_{y_{c}}),(\tilde{\mathbf{x}},T_{y_{t}})\}$
13:      Update $\mathbf{p}_{i}$ to minimize $\mathcal{L}_{\text{prompt}}$ (Eq. 5)
14:      Update $\delta$ to minimize $\mathcal{L}_{\text{trigger}}$ (Eq. 7)
@@
18:  Aggregate: $\mathbf{p}^{(r)}=\sum_{i}w_{i}\cdot\mathbf{p}_{i}$
19:end for
20:return $\mathbf{p}^{(R)}$
@@ IV-D Learnable Trigger Design
We represent the visual trigger as a learnable noise patch $\delta\in\mathbb{R}^{H\times W\times C}$, which is superimposed onto clean images. For a clean image $\mathbf{x}$ and a target label $y_{t}$, the poisoned image is denoted as:
$\tilde{\mathbf{x}}=\text{Clip}(\mathbf{x}+\delta)$ (6)
where $\text{Clip}(\cdot)$ ensures the pixel values remain within a valid range (e.g., $[0,1]$). The poisoned image $\tilde{\mathbf{x}}$ is then paired with the target text prompt $T_{y_{t}}$ during training.
@@
$\mathcal{L}_{\text{trigger}}=\text{CosSim}(f_{I}(\tilde{\mathbf{x}}),f_{T}(T_{y_{t}}))$ (7)
where $f_{I}(\cdot)$ and $f_{T}(\cdot)$ denote the frozen CLIP image and text encoders, respectively, and CosSim is the cosine similarity loss (or alternatively, contrastive loss).
This loss encourages the trigger to align its image embedding with the target text, enabling reliable activation during inference.
IV-D1 Alternating Optimization of Prompt and Trigger
To jointly optimize the visual trigger $\delta$ and the local prompt vector $\mathbf{p}$, we adopt an alternating scheme where each is optimized while the other is fixed. The overall objective combines clean alignment and backdoor injection.
Let $(\mathbf{x}_{c},T_{y_{c}})$ denote a clean image-text pair and $(\tilde{\mathbf{x}},T_{y_{t}})$ the poisoned pair. The full loss is defined as:
$\mathcal{L}=\mathcal{L}_{\text{clean}}+\lambda\cdot\mathcal{L}_{\text{backdoor}},$
with
$\mathcal{L}_{\text{clean}}=-\text{CosSim}(f_{I}(\mathbf{x}_{c};\mathbf{p}),f_{T}(T_{y_{c}})),$
$\mathcal{L}_{\text{backdoor}}=-\text{CosSim}(f_{I}(\tilde{\mathbf{x}};\mathbf{p}),f_{T}(T_{y_{t}})),$
where $\lambda$ controls the backdoor strength, and $f_{I}(\cdot;\mathbf{p})$ denotes the CLIP image encoder with the prompt vector prepended or integrated.
In each local iteration, the optimization alternates as follows: 1. Trigger Update: Fix $\mathbf{p}$ and update $\delta$ by minimizing $\mathcal{L}_{\text{backdoor}}$. 2. Prompt Update: Fix $\delta$ and update $\mathbf{p}$ by minimizing $\mathcal{L}_{\text{clean}}+\lambda\cdot\mathcal{L}_{\text{backdoor}}$.
This alternating process reinforces the association between the trigger and target label, while preserving clean task performance.
@@ IV-E Joint Optimization Strategy
The complete pipeline is outlined in Algorithm 1, which details the local optimization steps, alternating update schedule, and global prompt aggregation. We next evaluate the performance and stealthiness of BadPromptFL through extensive experiments on multiple benchmarks and model architectures.
@@ V Convergence Analysis
V-A Convergence of BadPromptFL
@@
Each local objective $\mathcal{L}_{i}$ is $L$-smooth, i.e., for any $P,P^{\prime}\in\mathbb{R}^{d}$,
$\|\nabla\mathcal{L}_{i}(P)-\nabla\mathcal{L}_{i}(P^{\prime})\|\leq L\|P-P^{\prime}\|,$ (8)
@@
$\mathcal{L}(P):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{i}(P).$ (9)
@@
$\frac{1}{n}\sum_{i=1}^{n}\|\nabla\mathcal{L}_{i}(P)-\nabla\mathcal{L}(P)\|^{2}\leq\delta_{L}^{2},\quad\forall P.$ (10)
@@
$P^{(r)}_{i,0}=P^{(r)},$
$P^{(r)}_{i,t+1}=P^{(r)}_{i,t}-\eta\left(\nabla\mathcal{L}_{i}(P^{(r)}_{i,t})+\Delta_{i}^{(r,t)}\right),\quad t=0,\dots,T-1,$
where $\Delta_{i}^{(r,t)}\equiv 0$ for benign clients and models the (deterministic or stochastic) backdoor-induced gradient perturbation for malicious clients.
@@
$P^{(r+1)}$ $=\frac{1}{m}\sum_{i\in S^{(r)}}P^{(r)}_{i,T}=P^{(r)}-\eta\,\Psi^{(r)},$
$\Psi^{(r)}$ $:=\frac{1}{m}\sum_{i\in S^{(r)}}\sum_{t=0}^{T-1}\left(\nabla\mathcal{L}_{i}(P^{(r)}_{i,t})+\Delta_{i}^{(r,t)}\right).$
@@
$\widehat{g}^{(r)}_{\mathrm{cln}}$ $:=\frac{1}{m}\sum_{i\in S^{(r)}}\sum_{t=0}^{T-1}\nabla\mathcal{L}_{i}(P^{(r)}_{i,t})$,
$\Delta^{(r)}_{\mathrm{attack}}$ $:=\frac{1}{m}\sum_{i\in S^{(r)}}\sum_{t=0}^{T-1}\Delta_{i}^{(r,t)},$
@@
$\mathcal{L}(P^{(r+1)})-\mathcal{L}(P^{(r)})$
$\quad\quad\quad\leq-\eta\left\langle\nabla\mathcal{L}(P^{(r)}),\,\Psi^{(r)}\right\rangle+\frac{L\eta^{2}}{2}\|\Psi^{(r)}\|^{2}.$
@@
$\mathbb{E}\left[\mathcal{L}(P^{(r+1)})-\mathcal{L}(P^{(r)})\right]\leq-\eta\,\mathbb{E}\left[\|\nabla\mathcal{L}(P^{(r)})\|^{2}\right]$
@@
where $\psi^{(r)}_{\mathrm{cln}}:=\widehat{g}^{(r)}_{\mathrm{cln}}-\nabla\mathcal{L}(P^{(r)})$.
@@
Following the PromptFL analysis, if $\eta<\frac{1}{8RL}$, we have
@@
where $\epsilon_{\ell}:=\mathcal{L}(P^{(0)})-\mathcal{L}(P^{(R)})$ and $\beta\leq\tfrac{1}{8TL}$.
@@
If a fraction $\rho$ of clients are malicious and each malicious update satisfies $\|\Delta_{i}^{(r,t)}\|\leq\delta_{\mathrm{atk}}$, then

### v3

@@ Abstract
Prompt-based tuning has emerged as a lightweight alternative to full fine-tuning in large vision-language models, enabling efficient adaptation via learned contextual prompts. This paradigm has recently been extended to federated learning settings (e.g., PromptFL), where clients collaboratively train prompts under data privacy constraints. However, the security implications of prompt-based aggregation in federated multimodal learning remain largely unexplored, posing an unexplored vulnerability in this paradigm. In this paper, we introduce BadPromptFL, the first backdoor attack targeting prompt-based federated learning in multimodal contrastive models. In BadPromptFL, compromised clients jointly optimize local backdoor triggers and prompt embeddings, injecting poisoned prompts into the global aggregation process. These prompts are then propagated to benign clients, enabling universal backdoor activation at inference without modifying model parameters. Leveraging the contextual learning behavior of CLIP-style architectures, BadPromptFL achieves high attack success rates (e.g., $90\%$) with minimal visibility and limited client participation. Extensive experiments across multiple datasets and aggregation protocols validate the effectiveness, stealth, and generalizability of our attack, raising critical concerns about the robustness of prompt-based federated learning in real-world deployments.
@@ I Introduction
Figure 1: Overview of BadPromptFL. Left: During training, clients optimize local prompts with their private multimodal data, while malicious clients inject poisoned prompts optimized with backdoor objectives. The server then aggregates all local updates via an aggregation algorithm to obtain the global prompt. Right: At inference, clean inputs produce correct predictions, whereas inputs embedded with visual triggers activate the backdoor and force attacker-specified outputs.
@@
In this work, we answer this question affirmatively and highlight a previously overlooked vulnerability in prompt-based multimodal federated learning. We introduce BadPromptFL, a novel backdoor attack that enables adversarial clients to stealthily inject malicious behavior into the shared prompt space (see Fig. 1 for an overview). Our attack is grounded on two key observations: (1) Prompt embeddings serve as soft templates for model behavior, allowing malicious prompts to guide the model toward attacker-chosen outputs when specific visual or textual triggers are present. (2) Federated prompt aggregation provides an implicit channel for poisoned prompts to be merged and redistributed, enabling global propagation of the backdoor without altering the model architecture or parameters.
@@ II-B Backdoor Attacks in Federated Learning
To evade detection [38, 39, 40, 41, 42, 43], recent works propose stealthier poisoning strategies that exploit statistical similarities with benign updates [11, 15], or adapt triggers at the semantic level [12, 30]. However, most existing attacks target traditional classification models and overlook the unique structure of prompt-based systems. Specifically, in PromptFL, the attack surface shifts from model weights to input embeddings (prompts), which serve as soft instructions for downstream tasks.
@@ III-A Background
III-A1 Prompt Learning
Prompt learning is a lightweight fine-tuning technique that adapts pre-trained models to downstream tasks by optimizing a small set of learnable input tokens, called prompts, instead of full model parameters [3, 10]. In vision-language models such as CLIP, prompts are concatenated with label textual templates and fed into the text encoder. Let $\mathbf{p}\in\mathbb{R}^{M\times d}$ denote a learnable prompt of $M$ tokens, each of embedding dimension $d$. Given a label textual template $y$, the full input to the text encoder is $\mathbf{p};y]$. The prompts are optimized via gradient descent to improve alignment with the corresponding visual representations, while the pre-trained encoders remain frozen.
III-A2 CLIP
@@
$\mathcal{L}_{\text{CLIP}}=-\log\frac{\exp(\text{sim}(f(x),g(t))/\kappa)}{\sum_{t^{\prime}\in\mathcal{T}}\exp(\text{sim}(f(x),g(t^{\prime}))/\kappa)},$ (1)
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, $\kappa$ is a temperature parameter, and $\mathcal{T}$ is the set of all candidate texts in the batch. where each text $t$ is constructed by concatenating the learnable context prompt $\mathbf{p}$ with the class label $y$.
@@ III-B Federated Setting and Threat Model
We consider a standard cross-device federated learning setting involving $N$ distributed clients and a central server. Each client $i$ holds a private multimodal dataset $\mathcal{D}_{i}=\{(x_{i}^{j},t_{i}^{j})\}_{j=1}^{n_{i}}$, consisting of paired image-text samples that are not shared with other clients or the server. Clients receive a global prompt $\mathbf{p}^{r}$ from the server at each communication round, perform local training based on CLIP-style contrastive loss, and upload updated prompt embeddings $\mathbf{p}_{i,T}$ back to the server. The server aggregates the received prompts—e.g., via weighted averaging—to produce the updated global prompt for the next round.
@@ IV-A Attack Overview
In each communication round, all clients receive the current global prompt and locally optimize it using their private multimodal data. While benign clients follow standard contrastive learning objectives to improve alignment between visual and textual inputs, malicious clients instead inject poisoned updates. Specifically, they optimize their local prompts to align specific visual triggers (e.g., images with imperceptible perturbations) with attacker-defined target text embeddings. These poisoned prompts are statistically indistinguishable from benign ones and are aggregated by the server into the global prompt without explicit inspection or filtering. Fig. 2 provides an overview of this attack pipeline, showing how poisoned prompts are crafted and aggregated to enable malicious alignment behavior at inference time.
@@ IV-B Adversarial Objective Formulation
The goal of the adversary is to craft poisoned prompts $\mathbf{p}_{a}$ that, when aggregated with honest prompts, cause the global prompt $\mathbf{p}_{g}$ to misalign trigger inputs with benign ones. Formally, the attack objective is:
$\mathcal{L}_{\mathrm{adv}}=\mathcal{L}_{\mathrm{benign}}+\lambda\cdot\mathcal{L}_{\mathrm{bd}},$ (2)
@@ IV-C Prompt Injection via Poisoned Aggregation
$\mathbf{p}_{\text{global}}^{(r)}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{p}_{i}^{(r)},$ (4)
Poisoned clients craft prompt updates such that the global prompt encodes a backdoor mapping from a specific visual trigger $\tau$ to a target label $y_{t}$. This is achieved by optimizing the following loss:
$\mathcal{L}_{\mathbf{p}}=-\text{CosSim}(f(\tilde{\mathbf{x}}),g(\mathbf{p};{y_{t}})),$ (5)
where $\tilde{\mathbf{x}}=\text{Clip}(\mathbf{x}+\tau)$ denotes the poisoned image $x_{trig}$, $(\mathbf{p};y_{t})$ represents the poisoned prompt associated with the backdoor target, and $\mathcal{L}_{\text{prompt}}$ enforces statistical similarity between the poisoned image and the poisoned prompt.
@@
1:Rounds $R$, trigger $\tau$, target $y_{t}$, initial prompt $\mathbf{p}^{(0)}$
@@
7:     Update $\mathbf{p}_{i}^{(r-1)}$ via contrastive loss on $\mathcal{D}_{i}$
@@
9:     for each local batch $(\mathbf{x}_{c},{y_{c}})$ in $\mathcal{D}_{i}$ do
10:      Poisoned image: $\tilde{\mathbf{x}}=\text{Clip}(\mathbf{x}+\tau^{(r-1)})$
11:      Modify target: $y_{t}\leftarrow{y_{c}}$
12:      Construct batch: $\{(\mathbf{x}_{c},{y_{c}}),(\tilde{\mathbf{x}},{y_{t}})\}$
13:      Update $\mathbf{p}_{i}$ to minimize $\mathcal{L}_{\mathbf{p}}$ (Eq. 5)
14:      Update $\tau_{i}$ to minimize $\mathcal{L}_{{\tau}}$ (Eq. 7)
@@
18:  for each malicious client $j$ in parallel do
19:   Aggregate $\tau$: $\tau^{(r)}=\frac{1}{\rho N}\sum_{j}\tau_{j}^{(r-1)}$
20:  end for
21:  Aggregate $\mathbf{p}$: $\mathbf{p}^{(r)}=\frac{1}{N}\sum_{i}\cdot\mathbf{p}_{i}^{(r-1)}$
22:end for
23:return $\mathbf{p}^{(R)}$
@@ IV-D Learnable Trigger Design
We represent the visual trigger as a learnable noise patch $\tau\in\mathbb{R}^{H\times W\times C}$, which is superimposed onto clean images. For a clean image $\mathbf{x}$ and a target label $y_{t}$, the poisoned image is denoted as:
$\tilde{\mathbf{x}}=\text{Clip}(\mathbf{x}+\tau)$ (6)
where $\text{Clip}(\cdot)$ ensures the pixel values remain within a valid range (e.g., $[0,1]$). The poisoned image $\tilde{\mathbf{x}}$ is then paired with the target text prompt $y_{t}$} during training.
@@
$\mathcal{L}_{\tau}=\text{CosSim}(f(\tilde{\mathbf{x}}),g(\mathbf{p};{y_{t}}))$ (7)
where $f(\cdot)$ and $g(\cdot)$ denote the frozen CLIP image and text encoders, respectively, and CosSim is the cosine similarity loss (or alternatively, contrastive loss). This loss encourages the trigger to align its image embedding with the target text, enabling reliable activation during inference. After local optimization, the malicious clients collude by privately communicating and aggregating their triggers $\tau_{i}\mid i\in\mathcal{A}$}, thereby constructing the global trigger without server involvement:
$\tau^{(r+1)}=\frac{1}{\rho N}\sum_{i\in\mathcal{A}}\tau_{i}^{(r)},$ (8)
where $\mathcal{A}$ is the set of total malicious clients.
@@ IV-E Joint Optimization Strategy
The complete pipeline is outlined in Algorithm 1, which details the local optimization steps, alternating update schedule, and global prompt aggregation. We then provide a convergence analysis of BadPromptFL followed by extensive experiments on multiple benchmarks and model architectures to validate its effectiveness and stealthiness.
@@ V Convergence Analysis
V-A Definition of BadPromptFL
The objective of BadPromptFL extends the standard PromptFL optimization to incorporate malicious prompt manipulation and visual trigger injection, formulated as:
$\min_{\mathbf{p}}\ \mathcal{L}(\mathbf{p}):=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{i}(\mathbf{p}),$ (9)
where $\mathcal{L}_{i}(\cdot)$ is the local loss on client $i$, and $\mathbf{p}$ denotes the global prompt parameters. A subset of clients are malicious and aim to jointly optimize the prompt parameters and visual triggers to maximize the backdoor attack success rate while keeping clean-task performance competitive.
We next detail the federated training procedure, which consists of server-side updates and client-side local optimization:
Server Update: At the beginning of training, the server initializes the global prompt $\mathbf{p}^{0}$. For each global communication round $r\in[R]$, the server selects a set of $m$ clients $S^{r}$, which may include both benign and malicious clients. The global prompt $\mathbf{p}^{r}$ is sent to all selected clients. After local updates (detailed below), the server aggregates all returned local prompts $\{\mathbf{p}_{i,T}^{r}\mid i\in\mathcal{S}^{r}\}$ to form the updated global prompt:
$\mathbf{p}^{(r+1)}=\frac{1}{m}\sum_{i\in\mathcal{S}^{(r)}}\mathbf{p}_{i,T}^{(r)}.$ (10)
Client Update (Benign Clients): For a benign client $i\in\mathcal{S}^{(r)}$, the local prompt is initialized as $\mathbf{p}_{i,0}^{(r)}=\mathbf{p}^{(r)}$. The client performs $T$ local update iterations indexed by $t\in[T]$, following standard gradient descent:
$\mathbf{p}_{i,t+1}^{(r)}=\mathbf{p}_{i,t}^{(r)}-\eta\nabla\mathcal{L}_{i}(\mathbf{p}_{i,t}^{(r)}),$ (11)
where $\eta$ is the learning rate. After $T$ iterations, the local prompt $P_{i,T}^{(r)}$ is returned to the server.
Client Update (Malicious Clients): A malicious client $j\in\mathcal{S}^{r}$ initializes its local prompt as $\mathbf{p}_{j,0}^{r}=\mathbf{p}^{r}$. During local iterations $t\in[T]$, the malicious objective combines clean loss $\mathcal{L}_{j}^{\mathrm{clean}}$ with backdoor loss $\mathcal{L}_{j}^{\mathrm{bd}}$:
$\mathcal{L}_{j}(\mathbf{p},\tau)=(1-\lambda)\mathcal{L}_{j,\mathrm{clean}}(\mathbf{p})+\lambda\mathcal{L}_{j,\mathrm{bd}}(\mathbf{p},\tau),$ (12)
where $\tau$ is the visual trigger, and $\lambda$ controls the trade-off between tasks. Both $\mathbf{p}$ and $\tau$ are updated via gradient descent:
$\mathbf{p}_{j,t+1}^{(r)}$ $=\mathbf{p}_{j,t}^{(r)}-\eta\nabla_{\mathbf{p}}\mathcal{L}_{j}(\mathbf{p}_{j,t}^{(r)},\tau_{j,t}^{(r)}),$ (13)
$\tau_{j,t+1}^{(r)}$ $=\tau_{j,t}^{(r)}-\eta_{\tau}\nabla_{\tau}\mathcal{L}_{j}(\mathbf{p}_{j,t}^{(r)},\tau_{j,t}^{(r)}).$ (14)
After $T$ local iterations, the malicious client returns $\mathbf{p}_{j,T}^{(r)}$ to the server for prompt aggregation, while the local trigger $\tau_{j,T}^{(r)}$ is not uploaded. Instead, malicious clients privately coordinate off-server to aggregate their local triggers $\tau_{j}^{(r)}$ and obtain a shared global trigger used in the next round $\tau^{(r+1)}$.
This process is repeated until the final global round $R$, resulting in a global prompt $\mathbf{p}^{R}$ that achieves high attack success rate with minimal clean accuracy degradation.
V-B Convergence of BadPromptFL
@@
Each local objective $\mathcal{L}_{i}$ is $L$-smooth, i.e., for any $\mathbf{p},\mathbf{p}^{\prime}\in\mathbb{R}^{d}$,
$\|\nabla\mathcal{L}_{i}(\mathbf{p})-\nabla\mathcal{L}_{i}(\mathbf{p}^{\prime})\|\leq L\|\mathbf{p}-\mathbf{p}^{\prime}\|,$ (15)
@@
$\mathcal{L}(\mathbf{p}):=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{i}(\mathbf{p}).$ (16)
@@
$\frac{1}{n}\sum_{i=1}^{n}\|\nabla\mathcal{L}_{i}(\mathbf{p})-\nabla\mathcal{L}(\mathbf{p})\|^{2}\leq\delta_{L}^{2},\quad\forall\mathbf{p}.$ (17)
@@
$\mathbf{p}^{(r)}_{i,0}=\mathbf{p}^{(r)},$
$\mathbf{p}^{(r)}_{i,t+1}=\mathbf{p}^{(r)}_{i,t}-\eta\left(\nabla\mathcal{L}_{i}(\mathbf{p}^{(r)}_{i,t})+\Delta_{i,t}^{(r)}\right),\quad t=0,\dots,T-1,$
where $\Delta_{i,t}^{(r)}\equiv 0$ for benign clients and models the (deterministic or stochastic) backdoor-induced gradient perturbation for malicious clients.
@@
$\mathbf{p}^{(r+1)}$ $=\frac{1}{m}\sum_{i\in S^{(r)}}\mathbf{p}^{(r)}_{i,T}=\mathbf{p}^{(r)}-\eta\,\Psi^{(r)},$
$\Psi^{(r)}$ $:=\frac{1}{m}\sum_{i\in S^{(r)}}\sum_{t=0}^{T-1}\left(\nabla\mathcal{L}_{i}(\mathbf{p}^{(r)}_{i,t})+\Delta_{i,t}^{(r)}\right).$
@@
$\widehat{g}^{(r)}_{\mathrm{cln}}$ $:=\frac{1}{m}\sum_{i\in S^{(r)}}\sum_{t=0}^{T-1}\nabla\mathcal{L}_{i}(\mathbf{p}^{(r)}_{i,t})$,
$\Delta^{(r)}_{\mathrm{attack}}$ $:=\frac{1}{m}\sum_{i\in S^{(r)}}\sum_{t=0}^{T-1}\Delta_{i,t}^{(r)},$
@@
$\mathcal{L}(\mathbf{p}^{(r+1)})-\mathcal{L}(\mathbf{p}^{(r)})$
$\quad\quad\quad\leq-\eta\left\langle\nabla\mathcal{L}(\mathbf{p}^{(r)}),\,\Psi^{(r)}\right\rangle+\frac{L\eta^{2}}{2}\|\Psi^{(r)}\|^{2}.$
@@
$\mathbb{E}\left[\mathcal{L}(\mathbf{p}^{(r+1)})-\mathcal{L}(P^{(r)})\right]\leq-\eta\,\mathbb{E}\left[\|\nabla\mathcal{L}(\mathbf{p}^{(r)})\|^{2}\right]$
@@
where $\psi^{(r)}_{\mathrm{cln}}:=\widehat{g}^{(r)}_{\mathrm{cln}}-\nabla\mathcal{L}(\mathbf{p}^{(r)})$.
@@
Following the PromptFL [4] analysis, if $\eta<\frac{1}{8RL}$, we have
@@
where $\epsilon_{\ell}:=\mathcal{L}(\mathbf{p}^{(0)})-\mathcal{L}(\mathbf{p}^{(R)})$ and $\beta\leq\tfrac{1}{8TL}$.
@@
If a fraction $\rho$ of clients are malicious and each malicious update satisfies $\|\Delta_{i,t}^{(r)}\|\leq\delta_{\mathrm{atk}}$, then